OpenStudy (anonymous):

A set of data items is normally distributed with a mean of 600 and a standard deviation of 20. Find the data item in this distribution that corresponds to the following z-score: z = 1.5

OpenStudy (kropot72):

\[z=\frac{X-\mu}{\sigma}\] \[1.5=\frac{X-600}{20}\] Now you can solve to find the data item X.

OpenStudy (anonymous):

I am still confused I do not know how to solve that equation, wil you please explain?

OpenStudy (kropot72):

Multiplying both sides by 20 gives: 30 = X - 600 X = 600 + 30 = 630

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)